# Learning to See and Act: Task-Aware View Planning for Robotic Manipulation

### v1

@@ Introduction
Table 3: Performance under different view numbers $K$ and camera radial constraints $r$, which $\Delta(r_{min},r_{max})=\pm 20\%$. Additionally, task PID $\to$ Put item in drawer, $T$ $\to$ Turn tap, PGC $\to$ Put groceries in cupboard, PMS $\to$ Put money in safe, CJ $\to$ Close jar.
Ablation Studies.
To verify the effectiveness of each module in our TAVP, we conducted comprehensive ablation studies on 18 RLBench tasks, as shown in Table 2. The full TAVP configuration achieves an 86.44% average success rate across all tasks. Removing the TaskMoE architecture from both feature extraction and action generation components (w/o TaskMoE) causes a moderate performance degradation to 85.56%, indicating its importance in multi-task representation learning. Crucially, replacing our actively explored viewpoints with random viewpoints (w/o Active Exploration) yields only 8.89% success, demonstrating that task-aware view planning contributes dominantly to the overall performance gains.
Method Average Success Rate (%) Average Inference Time(s)
ARP+ 84.90 0.394
TAVP 86.67 0.436
Table 4: Comparison of the average success rate and inference time between ARP+ and TAVP across 18 tasks.
TaskMoE PID $T$ PGC PMS CJ SR
Open drawer
(New task)
32.0 92.0 16.0 32.0 76.0 49.6 12.0
$\times$ 0.0 84.0 0.0 12.0 24.0 24.0 0.0
Table 5: Under the settings of $K=3$, $r_{\text{min}}=0.75$, and $r_{\text{max}}=1.3$, a comparison of performance on in-domain tasks (ID) and out-of-domain unseen tasks (OOD) with and without TaskMoE. Additionally, task PID $\to$ Put item in drawer, $T$ $\to$ Turn tap, PGC $\to$ Put groceries in cupboard, PMS $\to$ Put money in safe, CJ $\to$ Close jar.
Quantitative Analysis
To comprehensively evaluate our TAVP model, we conduct extensive experiments on five representative robotic manipulation tasks, as shown in Table 3. Our analysis encompasses three critical dimensions: 1) sensitivity analysis examining key model parameters including view quantity and camera radial constraints ($r_{min}$$r_{max}$), 2) performance-efficiency trade-off analysis quantifying inference latency against task success metrics, and 3) zero-shot generalization analysis of the TaskMoE component on unseen tasks.
Validation Analysis Experimental Setup. The validation analysis experiments were conducted on the RLBench dataset. We selected five manipulation tasks: put item in drawer, turn tap, put groceries in cupboard, put money in safe, and close jar. During training, each task utilized 20 demonstration episodes. For testing, each task was evaluated over 25 episodes, with the success rate recorded for each individual task and the average success rate calculated across all tasks. In the zero-shot generalization analysis experiment, the unseen task “Open drawer” was selected for evaluation. Similarly, this task was tested over 25 episodes. For the performance-efficiency trade-off analysis experiment, we maintained the setup described in Simulation Setup Simulation Setup..
Figure 4: Visualization of our TAVP and baseline model in simulation RLBench Environment and Real-world Environment.
Sensitivity Analysis.
We conducted ablation studies to analyze the key parameters of our TAVP model, including the camera radial constraint and the number of views. The radial constraint was evaluated in three cases based on a baseline range of ($0.75\sim 1.3$)m, which was varied by $\pm 20\%$. The experimental results, as shown in Table 3, indicate that compared to the baseline range, a smaller and tighter constraint ($0.90\sim 1.04$)m improved the model’s performance across all tasks, with the average success rate increasing significantly from $49.6\%$ to $56.0\%$. Conversely, a larger and looser constraint ($0.60\sim 1.56$)m led to a slight decrease in performance, with the average success rate dropping to $48.8\%$. This suggests that providing the model with a more precise and focused range of advantageous viewing distances can effectively enhance its final task performance.
In addition to radial constraints, we investigated the impact of varying the number of camera views on task performance. As shown in Table 3, under the same radial constraint range of $(0.75\sim 1.3)\text{m}$, increasing the number of views from 2 to 4 consistently improved performance across all five tasks. Specifically, the average success rate rose from $27.2\%$ (2 views) to $49.6\%$ (3 views), and further to $55.2\%$ with 4 views. Notably, performance gains were especially prominent in occlusion-prone tasks such as ”Put money in safe”, indicating that multi-view observations significantly enhance spatial perception and reduce failure caused by partial observability, enabling the TAVP to generalize better and make more informed decisions during manipulation tasks.
Performance-Efficiency Trade-off Analysis.
During inference, TAVP must predict dynamic multi-view camera poses and continuously adjust the rendering camera pose, resulting in a non-negligible computational overhead. To assess the practicality of the proposed TAVP model, we compare its average task success rate and inference latency with those of the baseline model, ARP+. The results, summarized in Table 4, show that TAVP achieves a higher average success rate than ARP+, while the increase in inference latency remains relatively modest (approximately 10.7%). These findings suggest that TAVP effectively balances task performance and operational efficiency, with design elements such as sampling acceleration and camera caching playing a key role in enhancing model efficiency.
Zero-shot Generalization Analysis.
We experimentally validated the generality and generalization capability of the TaskMoE module in the TAVP model, with the results presented in the Table 5. The experimental results indicate that, while keeping other parameters identical and at their default values, the inclusion of the TaskMoE module significantly enhances the model’s overall performance and generalization capability. On tasks seen during training, the model equipped with the TaskMoE module (✓) achieved an average success rate of 49.6%, more than double the 24.0% rate of the model without it ($\times$). This shows that the TaskMoE module enables more effective learning and execution of diverse, known tasks. For new, unseen tasks (OOD, Out-of-Distribution), the model without the TaskMoE module failed completely, with a success rate of 0. In contrast, the model with the TaskMoE module achieved a 12.0% success rate on the new task “Open drawer”, demonstrating a performance breakthrough. This proves that the TaskMoE endows the model with the ability to generalize learned knowledge to new tasks. In conclusion, the TaskMoE module not only enhances the model’s performance on known tasks but, more crucially, improves its generalization capability for new tasks, thereby validating the effectiveness of its design.
Visualization Results.
To validate the effectiveness of our method, we demonstrate the performance of TAVP and ARP+ in both the simulation environment RLBench and the real world. Figure 4 highlights TAVP’s advantage over the fixed-view ARP+ across simulation and reality. In RLBench, ARP+ loses sight of the mug once it is lifted, cannot perceive the cupboard interior, and drops the coffee mid-air; TAVP re-renders three task-guided views that simultaneously capture the mug and cupboard, yielding a correct 6-DoF placement. On the real robot, ARP+ misjudges the plate’s height, opening the gripper too early so grapes roll off target, whereas TAVP selects a clarifying side view and a persistent wrist view, preserving depth cues and completing the drop cleanly. By eliminating occlusion and recovering spatial context, TAVP turns visual completeness directly into manipulation success, validating that dynamic “seeing” underpins robust “acting.”
Conclusion
@@
Our TAVP framework establishes that dynamic view planning and task-aware representation learning significantly advance robotic manipulation capabilities. The MVEP module’s viewpoint optimization effectively overcomes occlusion limitations in fixed-viewpoint systems, while TaskMoE’s specialized feature extraction mitigates multi-task interference. These innovations collectively enhance performance across diverse manipulation challenges and enable meaningful generalization to novel tasks.
@@
In this supplementary material, we provide additional details and results of our proposed TAVP to complement the main paper. The content is organized as follows:
  • Appendix A: Real-world experimental details and results. Corresponding demo videos of the results are also included in the supplementary materials for a more intuitive understanding.
  • Appendix B: Additional design details of TAVP, including point cloud aggregation, camera pose parameterization, and the proposed pseudo-environment interaction mechanism.
  • Appendix C: Implementation details and information about the RLBench benchmark. Besides, we also provide more visualized results of RLBench [19], whose corresponding demo videos also included in the supplementary materials.
  • Appendix D: Discuss and analyze the limitations of our TAVP model.
Appendix A. Real-World Experimental Details and Results
Real-robot Setup.
We employ the Dobot Xtrainer platform featuring two 6-degree-of-freedom (6-DoF) Nova 2 collaborative manipulators, equipped with three Intel RealSense D405 depth cameras mounted at fixed positions: one at an overhead lateral position and two wrist-mounted cameras on both manipulators, as shown in Fig 5. Five distinct manipulation tasks were designed for evaluation the effectiveness of our proposed TAVP, including arrange_tableware, boil_water, cover_lid, pick_grape, and press_button. Expert demonstrations were collected, yielding 100 demonstration episodes for every task for training. Some samples are provided in Fig. 6. To accelerate the efficiency of training, we extract key frames from the collected demos and only retain important key frame data.
Real-world Experimental Results.
Fig. 7 shows the real-world experimental results. In five real-world tasks, our TAVP performs well in tasks such as boil_water, cover_lid, pick_grape, and press_button. Especially in the pick_grape task, it successfully executed despite occlusion of the manipulation target in one viewpoint, demonstrating the advantage of our model in global information perception. Corresponding demo videos are also provided for a more intuitive understanding of the real-world results.
@@ Appendix A. Real-World Experimental Details and Results
Figure 6: Demonstrations of 5 real-world tasks and their corresponding language instructions.
@@
Figure 7: Visualization of Robots’ Execution Processes in 5 Real-World Environments. In the figure, the T-axis represents the time axis, indicating the duration from the start to the end of the task. Each demonstration showcases views from two camera perspectives—the side camera and the wrist camera—along with the language instruction.
Appendix A Appendix B. More design details of TAVP
@@
In the RLBench dataset experiments, we select RGB-D images from four fixed viewpoints: front, right_shoulder, left_shoulder, and wrist as inputs. Following the approach of RVT-2 [143], we convert the depth maps from all viewpoints into point clouds in a global coordinate system using the intrinsic and extrinsic camera parameters. Subsequently, we perform multi-view point cloud aggregation and remove redundant points outside the workspace to obtain the final scene point cloud, as illustrated in Fig. 8.
@@
To enable efficient policy optimization without physical environment interaction, we design a novel pseudo-environment interaction mechanism. As shown in Algorithm 1. The shadow network $\pi_{\text{shadow}}$ (a frozen copy of the policy before training) processes observation $\mathbf{o}$ to compute reference losses $\mathcal{L}_{\text{ref}}$. MVEP $\pi_{\theta}$ processes $\mathbf{o}$ to generate camera poses $\mathbf{p}$ ($\mathbf{p}$ as policy action), the rendered observation images from multi-view camera poses are fed into TaskMoE and TaskMoE-ARP (Autoregressive Action Policy) to compute the current policy losses $\mathcal{L}_{\text{TAVP}}$. The reward $r$ is then calculated based on $\mathcal{L}_{\text{ref}}$, $\mathcal{L}_{\text{TAVP}}$, and $\mathbf{p}$. Finally, the experience tuple $(\mathbf{o},r,\mathbf{p},\log\pi_{\theta_{\text{old}}}(\mathbf{p}|\mathbf{o}),V_{\theta_{\text{old}}}(\mathbf{o}))$ is stored in the replay buffer $\mathcal{B}$ for policy updates. This approach enables batch environment interaction, leveraging existing demonstration data to improve policies and enhance data utilization efficiency.
@@
5:# Generate camera poses & compute current model loss using TAVP
@@
Camera Pose Parameterization
@@
Appendix B Appendix C. More Experimental Details and Resutls

### v4

@@ 4.2 Comparison with State-of-the-art Methods
Table 7: Performance under different view numbers $K$ and camera radial constraints $r$, which $\Delta(r_{min},r_{max})=\pm 20\%$. Task PID $\to$ Put item in drawer, $T$ $\to$ Turn tap, PGC $\to$ Put groceries in cupboard, PMS $\to$ Put money in safe, CJ $\to$ Close jar.
Setting of $K$ and $r$. We conducted ablation studies on the camera radial constraint $r$ and the number of views $K$. Compared to the baseline range ($0.75\sim 1.3$)m, a tighter constraint ($0.90\sim 1.04$)m improved performance, raising the average success rate from $49.6\%$ to $56.0\%$ (Table 7). A looser constraint ($0.60\sim 1.56$)m slightly reduced performance ($48.8\%$), indicating that a precise viewing distance range enhances task execution. Increasing the number of views from 2 to 4 under the baseline radial constraint consistently improved performance, with average success rates rising from $27.2\%$ to $49.6\%$ and $55.2\%$, respectively (Table 7). This demonstrates that multi-view observation mitigates occlusion and improves spatial perception. Considering computational cost, we chose $K=3$ as a balance.
5 Conclusion
@@
Our TVVE framework establishes that dynamic view planning and task-aware representation learning significantly advance robotic manipulation capabilities. The MVEP module’s viewpoint optimization effectively overcomes occlusion limitations in fixed-viewpoint systems, while TaskMoE’s specialized feature extraction mitigates multi-task interference. These contributions enhance performance across diverse manipulation challenges and enable meaningful generalization to novel tasks.
However, the framework has some limitations: multi-view re-rendering slightly increases inference latency, and reliance on accurate global point clouds makes it challenging to handle reflective or transparent objects. Future work could integrate multi-sensor data and domain adaptation to improve real-world robustness.
@@
In this supplementary material, we provide additional details and results of our proposed TVVE to complement the main paper. The content is organized as follows:
  • Appendix A: Real-world experimental details. Corresponding demo videos of the results are also included in the supplementary materials.
  • Appendix B: Additional design details of TVVE, including TaskMoE, point cloud aggregation, camera pose parameterization, and the proposed pseudo-environment interaction mechanism.
  • Appendix C: More simulation experimental details and results.
  • Appendix D: Implementation details and information about the RLBench [james2020rlbench] and RLBench-OG. Additionally, we provide more visualized results, with corresponding demo videos included in the supplementary materials.
  • Appendix E: Multi-view re-rendering visualization results.
Appendix A. Real-World Experimental Details
@@
To validate the generalizability of the proposed TVVE in real-world deployment scenarios, we conducted experiments on two distinct robotic platforms: the Dobot Nova 2 and the Franka Research 3, employing differentiated camera configurations (multi-view and single-view) to comprehensively validate the scenarios. For the Dobot Nova 2, a multi-view setup consisting of three perspectives (side, front, and wrist-mounted) was used, whereas for the Franka robot, only a front view was utilized. Experimental results confirm that our TVVE can be successfully deployed on both platforms and outperforms both Diffusion Policy and ARP. Figure 5 illustrates the real-world deployment environments and the corresponding images from each viewpoint.
Across five real-world tasks—Pick Grape, Stack Bowls, Push Buttons, Collect Fruits, and Put Item In Drawer—our TVVE achieves strong performance. Notably, in the pick_grape task, it successfully executed the manipulation even when the target was partially occluded in one viewpoint, demonstrating the advantage of our model in global perception capabilities. Corresponding demonstration videos are provided to facilitate a more intuitive understanding of the real-world outcomes.
Furthermore, to further evaluate the executive capability of our method across various types of tasks, we extended the experiments on the Franka Research 3 platform to include ten additional tasks. These tasks encompass articulated object manipulation, deformable object manipulation, among others. Experimental results are summarized in Table 8, which indicate that our TVVE achieves high success rates, highlighting its adaptability and effectiveness across a diverse set of tasks.
Arm Method Avg. SR (%) $\uparrow$ Pick Place Stack Bowls Push Buttons Collect Fruits Put Item In Drawer Rotate Handle Fold Towel Open Lip Unscrew Bottle Reach Drag
Franka ARP 62.0% 10/10 7/10 8/10 5/10 6/10 3/10 5/10 7/10 3/10 8/10
TVVE (Ours) 70.0% 10/10 7/10 9/10 6/10 7/10 3/10 5/10 9/10 5/10 9/10
Table 8: Task success rates for Franka robots across more manipulation tasks.
Real-World Robustness and Generalization Testing.
Arm Method Avg. SR (%) $\uparrow$ Seen Inst. Bkg. Obj. Occl. Illum.
Dobot Diffusion Policy 53.3 90.0 80.0 70.0 60.0 10.0 10.0
TVVE (Ours) 71.7 100.0 100.0 90.0 90.0 20.0 30.0
Table 9: Performance comparison under different configurations on the Dobot robot. Abbreviations: Inst. (Unseen Instance), Bkg. (Unseen Background), Obj. (Unseen Object), Occl. (Heavy Occlusion), Illum. (Illumination Variation).
We conduct robustness testing on the real-world Pick Grape task on the Dobot nova 2 robot, evaluating under various conditions including unseen instances, unseen backgrounds, unseen objects, heavy occlusion, and illumination variation. Throughout the experiments, heavy occlusion is identified as the primary cause of system failures. The results demonstrate that TVVE exhibits significantly better overall adaptability compared to the Diffusion Policy (DP) baseline. While both methods perform similarly in seen scenarios, TVVE substantially outperforms DP when confronted with unseen backgrounds, unseen objects, and especially under heavy occlusion. The success rates based on 10 trials for each configuration are reported, with quantitative comparisons detailed in Table 9.
@@
Figure 6: Visualization of Dobot and Franka Robots’ Execution Processes in Real-World Environments. In the figure, the T-axis represents the time axis, indicating the duration from the start to the end of the task. Each demonstration shows the following elements: side and wrist camera views for the Dobot robot, and a single front view for the Franka robot.
Details of Real-World Tasks.
We design real-world tasks to encompass as diverse operation types as possible, including grasping and placing, pressing/pushing, articulated object manipulation, rotation, deformable object manipulation, fine manipulation, tool usage, etc. The design details of each task are elaborated below.
Figure 7: Real-World Tasks Setup.
Pick and Place. On the table, there are fruits and a plate. The robot needs to pick up and place the specified fruits. Language instruction: Put the $\mathrm{[FRUIT]}$ into the $\mathrm{[COLOR]}$ plate. As shown in Fig. 7 (a)(f).
Stack Bowls. There are some bowls and a plate on the table. The robot needs to stack the bowls and plate in a specific order. Language instruction: Stack a $\mathrm{[COLOR]}$ bowl and a $\mathrm{[COLOR]}$ bowl on a $\mathrm{[COLOR]}$ plate. As shown in Fig. 7 (b)(g).
Push Buttons. There are multiple buttons of different colors on the table. The robot needs to press the buttons in a specific sequence. Language instruction: Press the $\mathrm{[COLOR]}$ button, then the $\mathrm{[COLOR]}$ button, followed by the $\mathrm{[COLOR]}$ button, and finally the $\mathrm{[COLOR]}$ button. As shown in Fig. 7 (c)(h).
Collect Fruits. There are various fruits on the table. The robot needs to collect the specified fruits into the specific target. Language instruction: Put a $\mathrm{[FRUIT]}$, a $\mathrm{[FRUIT]}$, and an $\mathrm{[FRUIT]}$ into the $\mathrm{[COLOR]}$ $\mathrm{[TARGET]}$. As shown in Fig. 7 (d)(i).
Put Item In Drawer. There is a cabinet on the table. The robot needs to open the specified drawer, place the target object inside, and then close the drawer. Language instruction: Open the $\mathrm{[Top/Middle/Bottom]}$ drawer and put the $\mathrm{[OBJECT]}$ into it. As shown in Fig. 7 (e)(j).
Rotate Handle. The experimenter holds the handle, and the robot needs to rotate the handle by a certain angle. Language instruction: Rotate the handle clockwise. As shown in Fig. 7 (k).
Fold Towel. There is a towel on the table. The robot needs to pick up one corner of the towel and then fold it. Language instruction: Fold the towel. As shown in Fig. 7 (l).
Open Lip. There is a box on the table. The robot needs to pick up the lid and open it. Language instruction: Open the plastic lid of the biscuit case. As shown in Fig. 7 (m).
Unscrew Bottle. There is a bottle on the table. The robot needs to twist the cap to open the bottle. Language instruction: Unscrew the black bottle. As shown in Fig. 7 (n).
Reach Drag. There is a tool and an object to be manipulated on the table. The robot needs to use the tool to move the target object to the target position. Language instruction: Sweep the green apple into the gap between buttons using wood board. As shown in Fig. 7 (o).
Appendix B. More design details of TVVE
TaskMoE: Mathematical Principles and Operational Mechanism
To handle heterogeneous manipulation tasks, we introduce a task-aware Mixture-of-Experts (TaskMoE) that routes visual tokens through (i) a task–instruction–aware gate layer and (ii) a per-gate expert selector. Unlike prior MoE work [shazeer2017outrageously, wangsparse], TaskMoE decouples tasks from gates, decouples gates from experts, and adds semantic/entropy regularizers to avoid collapse.
Cross-modal conditioning and FiLM. Let $\mathbf{V}\!\in\!\mathbb{R}^{S\times D}$ denote the sequence of visual tokens, where $S$ is the number of spatial tokens and $D$ is the feature dimension. Let $\mathbf{I}\!\in\!\mathbb{R}^{D}$ be the language instruction embedding and $\mathbf{t}\!\in\!\mathbb{R}^{D}$ the learnable embedding of the current task. We first fuse language into vision via cross-modal attention:
$\mathbf{Q}$ $=\mathbf{V}\mathbf{W}^{Q},\quad\mathbf{K}=\mathbf{I}\mathbf{W}^{K},\quad\mathbf{V}_{a}=\mathbf{I}\mathbf{W}^{V},$ (10)
$\mathbf{V}^{\prime}$ $=\mathrm{softmax}\!\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{a},$ (11)
where $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}\!\in\!\mathbb{R}^{D\times D}$ are projection matrices and $d_{k}$ is the key dimension. The attention output $\mathbf{V}^{\prime}\!\in\!\mathbb{R}^{S\times D}$ is an instruction-attended visual feature. We then modulate channels with FiLM using globally pooled vision and the task embedding (here $\mathrm{GAP}(\cdot)$ is global average pooling over the $S$ tokens):
$\mathbf{c}$ $=[\mathrm{GAP}(\mathbf{V}^{\prime});\,\mathbf{t}]\in\mathbb{R}^{2D},$ (12)
$\quad\gamma,\beta=\mathrm{MLP}(\mathbf{c})\in\mathbb{R}^{D}\times\mathbb{R}^{D},$ (13)
$\quad\mathbf{V}_{\text{mod}}=\gamma\odot\mathbf{V}+\beta,$ (14)
where $\gamma$ and $\beta$ are per-channel scaling and bias vectors and $\odot$ denotes element-wise multiplication. The modulated feature $\mathbf{V}_{\text{mod}}\!\in\!\mathbb{R}^{S\times D}$ serves as input to the gating network.
Stage 1: Task-aware gates ($N_{G}\!\ll\!N_{J}$). We consider $N_{J}$ tasks and $N_{G}$ latent gates. Gates represent reusable skill clusters shared across tasks. A shared gate head scores $N_{G}$ gates for each token and is modulated by a learnable task–gate assignment matrix $\mathbf{A}\in\mathbb{R}^{N_{J}\times N_{G}}$:
$\mathbf{p}_{g}$ $=\mathrm{softmax}(\mathbf{V}_{\text{mod}}\mathbf{W}_{g})\odot\mathrm{softmax}(\mathbf{A}[j,:]),$ (15)
where $\mathbf{W}_{g}\!\in\!\mathbb{R}^{D\times N_{G}}$ is the gate projection and $j\!\in\!\{1,\dots,N_{J}\}$ is the index of the current task. The matrix $\mathbf{p}_{g}\!\in\!\mathbb{R}^{S\times N_{G}}$ contains, for each token, a probability distribution over gates. We select the top-1 gate per token (keeping top-2 for analysis) and apply a per-gate capacity constraint to prevent any gate from absorbing too many tokens.
Stage 2: Per-gate expert selection ($N_{E}$ arbitrary). Given the chosen gate $g\!\in\!\{1,\dots,N_{G}\}$, we select experts from an independent pool of $N_{E}$ experts. Each gate has its own expert head:
$\mathbf{z}_{e}$ $=\mathbf{V}_{\text{mod}}\mathbf{W}_{\text{exp}}[g]+\mathbf{b}_{\text{exp}}[g],\quad\mathbf{p}_{e}=\mathrm{softmax}(\mathbf{z}_{e}),$ (16)
where $\mathbf{W}_{\text{exp}}\!\in\!\mathbb{R}^{D\times N_{G}\times N_{E}}$ and $\mathbf{b}_{\text{exp}}\!\in\!\mathbb{R}^{N_{G}\times N_{E}}$ are the gate–expert projection and bias, and $\mathbf{p}_{e}\!\in\!\mathbb{R}^{S\times N_{E}}$ gives, for each token, a probability distribution over experts under gate $g$. We keep top-$k$ experts per token (default k=2k\!=\!2), optionally filter the second expert with a threshold/random policy, and enforce an expert capacity across the sequence. The resulting dispatch tensor aggregates these expert weights to route tokens to experts, and expert outputs are combined using the selected mixture weights.
Regularization and loss. Let $\mathcal{L}_{\text{task}}$ denote the main imitation learning loss. We add balance, entropy, and semantic regularizers:
$\mathcal{L}$ $=\mathcal{L}_{\text{task}}+\lambda_{g}\mathcal{L}_{\text{bal-gate}}+\lambda_{e}\mathcal{L}_{\text{bal-exp}}$ (17)
$\quad-\eta_{g}H(\mathbf{p}_{g})-\eta_{e}H(\mathbf{p}_{e})+\lambda_{\text{sem}}\,\mathcal{L}_{\text{sem}}(\mathbf{A},\mathbf{I}),$ (18)
where $\mathcal{L}_{\text{bal-gate}}$ and $\mathcal{L}_{\text{bal-exp}}$ penalize unbalanced gate/expert utilization, $H(\mathbf{p}_{g})$ and $H(\mathbf{p}_{e})$ are the entropies of the gate and expert distributions (encouraging diverse routing), and $\mathcal{L}_{\text{sem}}(\mathbf{A},\mathbf{I})$ aligns rows of the task–gate matrix $\mathbf{A}$ with the similarity structure of instruction embeddings $\mathbf{I}$ so that semantically related tasks share gates/experts. The scalars $\lambda_{g},\lambda_{e},\eta_{g},\eta_{e},\lambda_{\text{sem}}\!\geq\!0$ control the strength of each regularizer. Capacity clipping at both stages further prevents any single gate or expert from monopolizing the traffic.
[…]
@@ Appendix B. More design details of TVVE
In the RLBench dataset experiments, we select RGB-D images from four fixed viewpoints: front, right_shoulder, left_shoulder, and wrist as inputs. Following the approach of RVT-2 [goyal2024rvt2], we convert the depth maps from all viewpoints into point clouds in a global coordinate system using the intrinsic and extrinsic camera parameters. Subsequently, we perform multi-view point cloud aggregation and remove redundant points outside the workspace to obtain the final scene point cloud, as illustrated in Fig. 8.
@@
To enable efficient policy optimization without physical environment interaction, we design a novel pseudo-environment interaction mechanism. As shown in Algorithm 1. The shadow network $\pi_{\text{shadow}}$ (a frozen copy of the policy before training) processes observation $\mathbf{o}$ to compute reference losses $\mathcal{L}_{\text{ref}}$. MVEP $\pi_{\theta}$ processes $\mathbf{o}$ to generate camera poses $\mathbf{p}$ ($\mathbf{p}$ as policy action), the rendered observation images from multi-view camera poses are fed into TaskMoE and TaskMoE-ARP (Autoregressive Action Policy) to compute the current policy losses $\mathcal{L}_{\text{TVVE}}$. The reward $r$ is then calculated based on $\mathcal{L}_{\text{ref}}$, $\mathcal{L}_{\text{TVVE}}$, and $\mathbf{p}$. Finally, the experience tuple $(\mathbf{o},r,\mathbf{p},\log\pi_{\theta_{\text{old}}}(\mathbf{p}|\mathbf{o}),V_{\theta_{\text{old}}}(\mathbf{o}))$ is stored in the replay buffer $\mathcal{B}$ for policy updates. This approach enables batch environment interaction, leveraging existing demonstration data to improve policies and enhance data utilization efficiency.
@@
5:# Generate camera poses & compute current model loss using TVVE
@@
Camera Pose Parameterization.
@@
Appendix C. More Experimental Details and Resutls
Baseline Details.
We compare with eleven state-of-the-art baselines on RLBench, RLbench-OG and Real-world, including both 2D and 3D approaches. For 2D methods, we include Diffusion Policy [chi2023diffusion], which formulates robot policies as conditional denoising diffusion processes. The 3D baselines comprise voxel-based methods: C2F-ARM-BC [james2022coarse], which predicts actions in a coarse-to-fine manner via Q-value estimation; PerAct [shridhar2023perceiver], which detects actions through global self-attention; Hiveformer [guhur2023instruction], which employs attention across historical features; and GNFactor [ze2023gnfactor], which co-optimizes a neural scene representation with a PerAct-based policy. We also consider point-based methods: PolarNet [chen2023polarnet], which computes dense point representations, and multi-view approaches: RVT [goyal2023rvt], which fuses multi-view predictions via 3D back-projection, and its successor RVT2 [goyal2024rvt2], which enhances precision and efficiency. Furthermore, we include Act3D [gervet2023act3d], which uses coarse-to-fine 3D featurization (retrained by us for a fair comparison), 3D Diffuser Actor [ke20253d], a 3D diffusion-based policy, and ARP [zhang2025autoregressive], an autoregressive policy generating hybrid action sequences via a Chunking Causal Transformer.